The formula R = l / (sigma S) is for thin wires.

(a) True

(b) False

(c) Not necessarily


1
Expert's answer
2021-07-21T09:51:16-0400

Gives

Thin wire

"R=\frac{\rho l}{S}\rightarrow(1)"

"\rho=\frac{1}{\sigma}\rightarrow(2)"

equation (1) and (2) we can written as

"R=\frac{ l}{\sigma S}\rightarrow(3)"

equation(3) is true

Option (a) correct